OpenStudy (anonymous):

33 sweatshirts were sold totaling $402.05. White ones sold for $11.95 and yellow for $12.50. How many of each were sold?

OpenStudy (anonymous):

a= number of whites, b= number of yellow, a+b=33, a*11.95 +b*12.50= 402.05, solve for either a or be and plug into the other: a= 33-b, (33-b)(11.95)+ b(12.50)= 402.05 plug b into a=33-b once b is known.

OpenStudy (anonymous):

i thougt b was 24 but it was wrong

OpenStudy (anonymous):

typo 23